OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition slope 4, and passes through (5,26).

OpenStudy (amorfide):

the equation you need is y-y1=m(x-x1) y1= y coordinate m=slope value x1= x coordinate

OpenStudy (anonymous):

^this is equivalent to y = mx+b

OpenStudy (anonymous):

which is also point slope formula

OpenStudy (anonymous):

i dont know how to set it up ??

OpenStudy (anonymous):

26 = 4(5) + b which is point slope formula 26 = 20 +b b = 6

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

no i dont... the question its self is confusing ?

OpenStudy (anonymous):

just think about look at the way i wrote it and solved everything, then you will understand